# Optimal rates of approximation by shallow ReLUk neural networks and applications to nonparametric regression

### 21 Jun 2023 v2

@@ 4.2 Deep neural networks and over-parameterization
There is a direct way to generalize the analysis in the last section to deep neural networks: we can implement shallow neural networks by sparse multi-layer neural networks with the same order of parameters, and estimate the approximation and generalization performance of the constructed networks. Since the optimal convergence rates of deep neural networks have already been established in [Schmidt-Hieber, 2020; Kohler and Langer, 2021], we do not pursue in this direction. Instead, we study the convergence rates of over-parameterized neural networks by using the idea discussed in [Jiao et al., 2023]. The reason for studying such networks is that, in modern applications of deep learning, the number of parameters in the networks is often much larger than the number of samples. However, in the convergence analysis of [Schmidt-Hieber, 2020; Kohler and Langer, 2021], the network that achieves the optimal rate is under-parameterized (see also the choice of $N_{n}$ in Theorem 4.2). Hence, these analysis can not explain the empirical performance of deep learning models used in practice.
Following Jiao et al. [2023], we consider deep neural networks with norm constraints on weight matrices. For $W,L\in\mathbb{N}$, we denote by $\mathcal{NN}(W,L)$ the set of functions that can be parameterized by ReLU neural networks in the form
@@
where we use $\|A\|:=\sup_{\|x\|_{\infty}\leq 1}\|Ax\|_{\infty}$ to denote the operator norm (induced by the $\ell^{\infty}$ norm) of a matrix $A=(a_{i,j})\in\mathbb{R}^{m\times n}$. It is well-known that $\|A\|$ is the maximum $1$-norm of the rows of $A$:
@@
The motivation for such a definition of $\kappa(\theta)$ is discussed in [Jiao et al., 2023]. For $M\geq 0$, we denote by $\mathcal{NN}(W,L,M)$ as the set of functions $f_{\theta}\in\mathcal{NN}$ that satisfy $\kappa(\theta)\leq M$. It is shown by [Jiao et al., 2023, Proposition 2.5] that, if $W_{1}\leq W_{2},L_{1}\leq L_{2},M_{1}\leq M_{2}$, then $\mathcal{NN}(W_{1},L_{1},M_{1})\subseteq\mathcal{NN}(W_{2},L_{2},M_{2})$. (Strictly speaking, Jiao et al. [2023] use the convention that the bias $b^{(L)}=0$ in the last layer. But the results can be easily generalized to the case $b^{(L)}\neq 0$, see [Yang, 2022, Section 2.1] for details.)
@@
On the contrary, let $f_{\theta}\in\mathcal{NN}(N,1,M)$ be a function parameterized in the form (4.3) with $(A^{(0)},b^{(0)})=(a^{(0)}_{1},\dots,a^{(0)}_{N})^{\intercal}$ and $(A^{(1)},b^{(1)})=(a^{(1)}_{1},\dots,a^{(1)}_{N},b^{(1)})$, where $a^{(0)}_{i}\in\mathbb{R}^{d+1}$ and $a^{(1)}_{i},b^{(1)}\in\mathbb{R}$. Then, $f_{\theta}$ can be represented as
@@
Part (1) is a direct consequence of Corollary 2.4 and $\mathcal{F}_{\sigma}(W,M)\subseteq\mathcal{NN}(W,L,\sqrt{d+1}M)$. Part (2) follows from Lemma 2.3. ∎
In the first part of Corollary 4.5, if we allow the width $W$ to be arbitrary large, say $W\gtrsim M^{2d/(d+3-2\alpha)}$, then we can bound the approximation error by the size of weights. Hence, this result can be applied to over-parameterized neural networks. (Note that, in Theorem 4.2, we use a different regime of the bound.) In part (2), the size of weights is bounded by a constant. We will show that this constant can be used to control the generalization error. Since the approximation error is bounded by $W$ and is independent of $M$, we do not have trade-off in the error decomposition of ERM and only need to choose $W$ sufficiently large to reduce the approximation error. Hence, it can also be applied to over-parameterized neural networks.
The approximation rate $M^{-\frac{2\alpha}{d+3-2\alpha}}$ for $\mathcal{H}^{\alpha}$ in Corollary 4.5 improves the rate $M^{-\frac{\alpha}{d+1}}$ proved by Jiao et al. [2023]. Using the upper bound for Rademacher complexity of $\mathcal{NN}(W,L,M)$ (see Lemma 4.6), Jiao et al. [2023] also gave an approximation lower bound $(M\sqrt{L})^{-\frac{2\alpha}{d-2\alpha}}$. For fixed depth $L$, our upper bound is very close to this lower bound. We conjecture that the rate in Corollary 4.5 is optimal with respect to $M$ (for fixed depth $L$). The discussion of optimality in the last part of Section 2 implies that the conjecture is true for shallow neural networks.
To control the generalization performance of over-parameterized neural networks, we need to have size-independent sample complexity bounds for such networks. Several methods have been applied to obtain such kind of bounds in recent works [Neyshabur et al., 2015, 2018; Bartlett et al., 2017; Golowich et al., 2020]. Here, we will use the result of Golowich et al. [2020], which estimates the Rademacher complexity of deep neural networks [Bartlett and Mendelson, 2002]. For a set $S\subseteq\mathbb{R}^{n}$, let us denote its Rademacher complexity by
@@
where $\xi_{1:n}=(\xi_{1},\dots,\xi_{n})$ is a sequence of i.i.d. Rademacher random variables. The following lemma is from Golowich et al. [2020, Theorem 3.2] and Jiao et al. [2023, Lemma 2.3].
Lemma 4.6 (Golowich et al. [2020]).
@@
Now, we can estimate the convergence rates of the ERM based on over-parameterized neural networks. As usual, we decompose the excess risk of the ERM into approximation error and generalization error, and bound them by Corollary 4.5 and Lemma 4.6, respectively. Note that the convergence rates in the following theorem are very close to the optimal rates in Theorem 4.2.
@@
Let $d\geq 4$ and $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\{\mathcal{T}_{B_{n}}f:f\in\mathcal{NN}(W_{n},L,M_{n})\}$, where $L\in\mathbb{N}$ is a fixed constant, $1\leq B_{n}\lesssim\log n$ in case (1) and $\sqrt{2}\leq B_{n}\lesssim\log n$ in case (2).
@@
    $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}},\quad W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}},$
@@
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then there exist a constant $M\geq 1$ such that for $M_{n}=M$ and any $W_{n}\gtrsim n^{d/(2d+6)}$, we have
    $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-1/2}\log n.$
@@
The proof is essentially the same as Jiao et al. [2023, Theorem 4.1]. Observe that, for any $f\in\mathcal{F}_{n}$,
@@
$=$ $\mathbb{E}_{X_{1:n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}-\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}\right]+2\mathbb{E}_{\mathcal{D}_{n}}\left[\frac{1}{n}\sum_{i=1}^{n}\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))\right]$
@@
where we denote $\Phi_{n}:=\{f-h:f\in\mathcal{F}_{n}\}$. By a standard symmetrization argument (see Wainwright [2019, Theorem 4.10]), we can bound the first term in (4.5) by the Rademacher complexity:
@@
where $\Phi_{n}^{2}(X_{1:n}):=\{(\phi^{2}(X_{1}),\dots,\phi^{2}(X_{n}))\in\mathbb{R}^{n}:\phi\in\Phi_{n}\}\subseteq\mathbb{R}^{n}$ is the set of function values on the sample points. Recall that we assume $B_{n}\geq 1$ in case (1) and $B_{n}\geq\sqrt{2}$ in case (2). Hence, we always have $\|h\|_{L^{\infty}(\mathbb{B}^{d})}\leq\sqrt{2}$ and $\|\phi\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2B_{n}$ for any $\phi\in\Phi_{n}$. By the structural properties of Rademacher complexity [Bartlett and Mendelson, 2002, Theorem 12],
@@
where we apply Lemma 4.6 in the last inequality. Note that the second term in (4.5) is a Gaussian complexity. We can also bound it by the Rademacher complexity [Bartlett and Mendelson, 2002, Lemma 4]:
@@
Combining with (4.4) and (4.6), if we choose $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}}$ and $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}}$, then
@@ 4.3 Convolutional neural networks
In contrast to the vast amount of theoretical studies on fully connected neural networks, there are only a few papers analyzing the performance of convolutional neural networks [Zhou, 2020a, b; Fang et al., 2020; Petersen and Voigtlaender, 2020; Lin et al., 2022]. The recent work [Lin et al., 2022] showed the universal consistency of CNNs for nonparametric regression. In this section, we show how to use our approximation results to analyze the convergence rates of CNNs.
Following [Zhou, 2020b], we introduce a sparse convolutional structure on deep neural networks. Let $s\geq 2$ be a fixed integer, which is used to control the filter length. Given a sequence $w=(w_{i})_{i\in\mathbb{Z}}$ on $\mathbb{Z}$ supported on $\{0,1,\dots,s\}$, the convolution of the filter $w$ with another sequence $x=(x_{i})_{i\in\mathbb{Z}}$ supported on $\{1,\dots,d\}$ is a sequence wxw*x given by
@@
with $N_{\ell}-2s$ repeated components in the middle. By definition, it is easy to see that $\mathcal{CNN}(s,L)\subseteq\mathcal{NN}(d+Ls,L)$. The assumption on the special form (4.7) of biases is used to reduce the free parameters in the network. As in [Zhou, 2020b], one can compute that the number of free parameters in $\mathcal{CNN}(s,L)$ is (5s+2)L+2d2s(5s+2)L+2d-2s, which grows linearly on $L$.
@@
This result is proved in Zhou [2020b, Proof of Theorem 2]. We only give a sketch of the construction for completeness. Any function $f\in\mathcal{NN}(N,1)$ can be written as $f(x)=\sum_{i=1}^{N}c_{i}\sigma(a_{i}^{\intercal}x+b_{i})+c_{0}$, where $a_{i}\in\mathbb{R}^{d}$ and $b_{i},c_{i}\in\mathbb{R}$. Define a sequence $v$ supported on $\{0,\dots,Nd-1\}$ by stacking the vectors $a_{1},\dots,a_{N}$ (with components reversed) by
@@
Applying Zhou [2020b, Theorem 3] to the sequence $v$, we can construct filters $\{w^{(\ell)}\}_{\ell=0}^{L-1}$ supported on $\{0,1,\dots,s\}$ such that $v=w^{(L-1)}*w^{(L-2)}*\dots*w^{(0)}$, which implies $A^{w^{(L-1)}}\cdots A^{w^{(0)}}=A^{v}\in\mathbb{R}^{(d+Ls)\times d}$. Note that, by definition, for $i=1,\dots,N$, the $id$-th row of $A^{v}$ is exactly $a_{i}^{\intercal}$. Then, for $\ell=0,\dots,L-2$, we can choose $b^{(\ell)}$ satisfying (4.7) such that $f^{(\ell+1)}(x)=A^{w^{(\ell)}}\cdots A^{w^{(0)}}x+B^{(\ell)}$, where $B^{(\ell)}>0$ is a sufficiently large constant that makes the components of $f^{(\ell+1)}(x)$ positive for all $x\in\mathbb{B}^{d}$. Finally, we can construct $b^{(L-1)}$ such that $f^{(L)}_{k}(x)=\sigma(a_{i}^{\intercal}x+b_{i})$ for $i=1,\dots,N$ and $k=id$, which implies $f\in\mathcal{CNN}(s,L)$. ∎
@@
Since the number of parameters in $\mathcal{CNN}(s,L)$ is approximately $L$, the rate $\mathcal{O}(L^{-\alpha/d})$ in part (1) of Corollary 4.9 is the same as the rate in [Yarotsky, 2017] for fully connected neural networks. However, Yarotsky [2018]; Lu et al. [2021] showed that this rate can be improved to $\mathcal{O}(L^{-2\alpha/d})$ for fully connected neural networks by using the bit extraction technique [Bartlett et al., 2019]. It would be interesting to see whether this rate also holds for $\mathcal{CNN}(s,L)$.
As in Theorem 4.2, we use Lemma 4.1 to decompose the error and bound the approximation error by Corollary 4.9. But we need to use pseudo-dimension to bound the covering number, because we do not have estimate for the size of weights in CNNs.
@@
Let $d\geq 4$ and $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{CNN}(s,L_{n})$, where $s\geq 2$ is a fixed integer, and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
@@
To apply Lemma 4.1, we need to estimate the covering number $\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})$. The classical result of Haussler [1992, Theorem 6] showed that the covering number can be bounded by pseudo-dimension:
$\log\mathcal{N}(\epsilon,\mathcal{T}_{B_{n}}\mathcal{F}_{n},\|\cdot\|_{L^{1}(X_{1:n})})\lesssim\,{\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})\log(B_{n}/\epsilon),$
where ${\rm Pdim}\,(\mathcal{T}_{B_{n}}\mathcal{F}_{n})$ is the pseudo-dimension of the function class $\mathcal{T}_{B_{n}}\mathcal{F}_{n}$, see (2.2). For ReLU neural networks, Bartlett et al. [2019] gave the following estimate of the pseudo-dimension:

### 09 Jan 2024 v3

@@ 4.2 Deep neural networks and over-parameterization
There is a direct way to generalize the analysis in the last section to deep neural networks: we can implement shallow neural networks by sparse multi-layer neural networks with the same order of parameters, and estimate the approximation and generalization performance of the constructed networks. Since the optimal convergence rates of deep neural networks have already been established in [52, 27], we do not pursue in this direction. Instead, we study the convergence rates of over-parameterized neural networks by using the idea discussed in [24]. The reason for studying such networks is that, in modern applications of deep learning, the number of parameters in the networks is often much larger than the number of samples. However, in the convergence analysis of [52, 27], the network that achieves the optimal rate is under-parameterized (see also the choice of $N_{n}$ in Theorem 4.2). Hence, the analysis can not explain the empirical performance of deep learning models used in practice.
Following [24], we consider deep neural networks with norm constraints on weight matrices. For $W,L\in\mathbb{N}$, we denote by $\mathcal{NN}(W,L)$ the set of functions that can be parameterized by ReLU neural networks in the form
@@
where we use $\|A\|:=\sup_{\|x\|_{\infty}\leq 1}\|Ax\|_{\infty}$ to denote the operator norm (induced by the $\ell^{\infty}$ norm) of a matrix $A=(a_{i,j})\in\mathbb{R}^{m\times n}$. It is well-known that $\|A\|$ is the maximum 11-norm of the rows of AA:
@@
The motivation for such a definition of $\kappa(\theta)$ is discussed in [24]. For $M\geq 0$, we denote by $\mathcal{NN}(W,L,M)$ as the set of functions $f_{\theta}\in\mathcal{NN}(W,L)$ that satisfy $\kappa(\theta)\leq M$. It is shown by [24, Proposition 2.5] that, if $W_{1}\leq W_{2},L_{1}\leq L_{2},M_{1}\leq M_{2}$, then $\mathcal{NN}(W_{1},L_{1},M_{1})\subseteq\mathcal{NN}(W_{2},L_{2},M_{2})$. (Strictly speaking, [24] use the convention that the bias $b^{(L)}=0$ in the last layer. But the results can be easily generalized to the case $b^{(L)}\neq 0$, see [62, Section 2.1] for details.)
@@
On the other side, let $f_{\theta}\in\mathcal{NN}(N,1,M)$ be a function parameterized in the form (4.4) with $(A^{(0)},b^{(0)})=(a^{(0)}_{1},\dots,a^{(0)}_{N})^{\intercal}$ and $(A^{(1)},b^{(1)})=(a^{(1)}_{1},\dots,a^{(1)}_{N},b^{(1)})$, where $a^{(0)}_{i}\in\mathbb{R}^{d+1}$ and $a^{(1)}_{i},b^{(1)}\in\mathbb{R}$. Then, $f_{\theta}$ can be represented as
@@
The first part is a direct consequence of Corollary 2.4 and the inclusion $\mathcal{F}_{\sigma}(W,M)\subseteq\mathcal{NN}(W,L,\sqrt{d+1}M)$. The second part follows from Lemma 2.3 and we can choose $M=\sqrt{d+1}$. ∎
In the first part of Corollary 4.5, if we allow the width $W$ to be arbitrary large, say $W\gtrsim M^{2d/(d+3-2\alpha)}$, then we can bound the approximation error by the size of weights. Hence, this result can be applied to over-parameterized neural networks. (Note that, in Theorem 4.2, we use a different regime of the bound.) For the approximation of $\mathcal{F}_{\sigma}(1)$, the size of weights is bounded by a constant. We will show that this constant can be used to control the generalization error. Since the approximation error is bounded by $W$ and is independent of $M$, we do not have trade-off in the error decomposition of ERM and only need to choose $W$ sufficiently large to reduce the approximation error. Hence, it can also be applied to over-parameterized neural networks.
The approximation rate $M^{-\frac{2\alpha}{d+3-2\alpha}}$ for $\mathcal{H}^{\alpha}$ in Corollary 4.5 improves the rate $M^{-\frac{\alpha}{d+1}}$ proven by [24]. Using the upper bound for Rademacher complexity of $\mathcal{NN}(W,L,M)$ (see Lemma 4.6), [24] also gave an approximation lower bound $(M\sqrt{L})^{-\frac{2\alpha}{d-2\alpha}}$. For fixed depth $L$, our upper bound is very close to this lower bound. We conjecture that the rate in Corollary 4.5 is optimal with respect to $M$ (for fixed depth $L$). The discussion of optimality at the end of Section 2 implies that the conjecture is true for shallow neural networks (i.e. $L=1$).
To control the generalization performance of over-parameterized neural networks, we need to have size-independent sample complexity bounds for such networks. Several methods have been applied to obtain such kind of bounds in recent works [43, 42, 5, 21]. Here, we will use the result of [21], which estimates the Rademacher complexity of deep neural networks [7]. For a set $S\subseteq\mathbb{R}^{n}$, let us denote its Rademacher complexity by
@@
where $\xi_{1:n}=(\xi_{1},\dots,\xi_{n})$ is a sequence of i.i.d. Rademacher random variables. The following lemma is from [21, Theorem 3.2] and [24, Lemma 2.3].
Lemma 4.6.
@@
Now, we can estimate the convergence rates of the ERM based on over-parameterized neural networks. As usual, we decompose the excess risk of the ERM into approximation error and generalization error, and bound them by Corollary 4.5 and Lemma 4.6, respectively. Note that the convergence rates in the following theorem are worse than the optimal rates in Theorem 4.2.
@@
Let $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\{\mathcal{T}_{B_{n}}f:f\in\mathcal{NN}(W_{n},L,M_{n})\}$, where $L\in\mathbb{N}$ is a fixed constant, $1\leq B_{n}\lesssim\log n$ in case (1) and $\sqrt{2}\leq B_{n}\lesssim\log n$ in case (2).
@@
    $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}},\quad M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}},$
@@
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we choose a large enough constant $M$ and let
    $W_{n}\gtrsim n^{\frac{d}{2d+6}},\quad M_{n}=M,$
    then
    $\mathbb{E}_{\mathcal{D}_{n}}\|f_{n}^{*}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{1}{2}}\log n.$
@@
The proof is essentially the same as [24, Theorem 4.1]. Observe that, for any $f\in\mathcal{F}_{n}$,
@@
$=$ $\mathbb{E}_{\mathcal{D}_{n}}\left[\|f_{n}^{*}-h\|^{2}_{L^{2}(\mu)}-\frac{1}{n}\sum_{i=1}^{n}(f_{n}^{*}(X_{i})-h(X_{i}))^{2}\right]+2\mathbb{E}_{\mathcal{D}_{n}}\left[\frac{1}{n}\sum_{i=1}^{n}\eta_{i}(f_{n}^{*}(X_{i})-h(X_{i}))\right]$
@@
where we denote $\Phi_{n}:=\{f-h:f\in\mathcal{F}_{n}\}$. By a standard symmetrization argument (see [60, Theorem 4.10]), we can bound the first term in (4.6) by the Rademacher complexity:
@@
where $\Phi_{n}^{2}(X_{1:n}):=\{(\phi^{2}(X_{1}),\dots,\phi^{2}(X_{n}))\in\mathbb{R}^{n}:\phi\in\Phi_{n}\}\subseteq\mathbb{R}^{n}$ is the set of function values on the sample points. Recall that we assume $B_{n}\geq 1$ in case (1) and $B_{n}\geq\sqrt{2}$ in case (2). Hence, we always have $\|h\|_{L^{\infty}(\mathbb{B}^{d})}\leq\sqrt{2}$ and $\|\phi\|_{L^{\infty}(\mathbb{B}^{d})}\leq 2B_{n}$ for any $\phi\in\Phi_{n}$. By the structural properties of Rademacher complexity [7, Theorem 12],
@@
where we apply Lemma 4.6 in the last inequality. Note that the second term in (4.6) is a Gaussian complexity. We can also bound it by the Rademacher complexity [7, Lemma 4]:
@@
Combining with (4.5) and (4.7), we know that if we choose $M_{n}\asymp n^{\frac{1}{2}-\frac{2\alpha}{d+3+2\alpha}}$ and $W_{n}\gtrsim n^{\frac{d}{d+3+2\alpha}}$, then
@@ 4.3 Convolutional neural networks
In contrast to the vast amount of theoretical studies on fully connected neural networks, there are only a few papers analyzing the performance of convolutional neural networks [66, 67, 16, 47, 30, 20, 35, 68]. The recent work [30] showed the universal consistency of CNNs for nonparametric regression. In this section, we show how to use our approximation results to analyze the convergence rates of CNNs.
Following [67], we introduce a sparse convolutional structure on deep neural networks. Let $s\geq 2$ be a fixed integer, which is used to control the filter length. Given a sequence $w=(w_{i})_{i\in\mathbb{Z}}$ on $\mathbb{Z}$ supported on $\{0,1,\dots,s\}$, the convolution of the filter $w$ with another sequence $x=(x_{i})_{i\in\mathbb{Z}}$ supported on $\{1,\dots,d\}$ is a sequence wxw*x given by
@@
with $N_{\ell}-2s$ repeated components in the middle. By definition, it is easy to see that $\mathcal{CNN}(s,L)\subseteq\mathcal{NN}(d+Ls,L)$. The assumption on the special form (4.8) of biases is used to reduce the free parameters in the network. As in [67], one can compute that the number of free parameters in $\mathcal{CNN}(s,L)$ is (5s+2)L+2d2s(5s+2)L+2d-2s, which grows linearly on $L$.
@@
This result is proven in [67, Proof of Theorem 2]. We only give a sketch of the construction for completeness. Any function $f\in\mathcal{NN}(N,1)$ can be written as $f(x)=\sum_{i=1}^{N}c_{i}\sigma(a_{i}^{\intercal}x+b_{i})+c_{0}$, where $a_{i}\in\mathbb{R}^{d}$ and $b_{i},c_{i}\in\mathbb{R}$. Define a sequence $v$ supported on $\{0,\dots,Nd-1\}$ by stacking the vectors $a_{1},\dots,a_{N}$ (with components reversed) by
@@
Applying [67, Theorem 3] to the sequence $v$, we can construct filters $\{w^{(\ell)}\}_{\ell=0}^{L-1}$ supported on $\{0,1,\dots,s\}$ such that $v=w^{(L-1)}*w^{(L-2)}*\dots*w^{(0)}$, which implies $A^{w^{(L-1)}}\cdots A^{w^{(0)}}=A^{v}\in\mathbb{R}^{(d+Ls)\times d}$. Note that, by definition, for $i=1,\dots,N$, the idid-th row of $A^{v}$ is exactly $a_{i}^{\intercal}$. Then, for $\ell=0,\dots,L-2$, we can choose $b^{(\ell)}$ satisfying (4.8) such that $f^{(\ell+1)}(x)=A^{w^{(\ell)}}\cdots A^{w^{(0)}}x+B^{(\ell)}$, where $B^{(\ell)}>0$ is a sufficiently large constant that makes the components of $f^{(\ell+1)}(x)$ positive for all $x\in\mathbb{B}^{d}$. Finally, we can construct $b^{(L-1)}$ such that $f^{(L)}_{k}(x)=\sigma(a_{i}^{\intercal}x+b_{i})$ for $i=1,\dots,N$ and $k=id$, which implies $f\in\mathcal{CNN}(s,L)$. ∎
@@
Since the number of parameters in $\mathcal{CNN}(s,L)$ is approximately $L$, the rate $\mathcal{O}(L^{-\alpha/d})$ in part (1) of Corollary 4.9 is the same as the rate in [64] for fully connected neural networks. However, [65, 31] showed that this rate can be improved to $\mathcal{O}(L^{-2\alpha/d})$ for fully connected neural networks by using the bit extraction technique [6]. It would be interesting to see whether this rate also holds for $\mathcal{CNN}(s,L)$.
As in Theorem 4.2, we use Lemma 4.1 to decompose the error and bound the approximation error by Corollary 4.9. The covering number is bounded again by pseudo-dimension.
@@
Let $f_{n}^{*}$ be the estimator (4.2) with $\mathcal{F}_{n}=\mathcal{CNN}(s,L_{n})$, where $s\geq 2$ is a fixed integer, and set $B_{n}=c_{1}\log n$ for some constant $c_{1}>0$.
@@
The proof is the same as Theorem 4.2 and [68]. We can use (4.3) to bound the covering number by the pseudo-dimension. For convolutional neural networks, [6] gave the following estimate of the pseudo-dimension:
@@
Finally, we note that the recent paper [68] also studied the convergence of CNNs and proved the rate $\mathcal{O}(n^{-1/3}(\log n)^{2})$ for $\mathcal{H}^{\alpha}$ with $\alpha>(d+4)/2$. The convergence rate we obtained in Theorem 4.10 for $\mathcal{F}_{\sigma}(1)$, which includes $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ by Theorem 2.1, is slightly better than their rate.